# The CN matrix of a pure braid projection

### v2

@@ 2 Properties of CN matrix
for some $11\leq I\leq n-m+$ 1. Then $M_{2}$ is CN-realizable.
@@
Let $B_{1}$ be an $m$-braid projection such that $N(B_{1})=M_{1}$. Place $I-1$ strands on the left-hand side and $n-m-I+1$ strands on the right-hand side to obtain an $n$-braid projection $B_{2}$. Then $N(B_{2})=M_{2}$. (See Figure 3.) ∎
@@
Let $M$ be a $6\times 6$ T0 strictly upper triangular $(0,2)$-matrix such that $M(i,j)=0$ when $i<I$ or $j>I-m-1$ for some integers $1\leq m\leq 5$ and $1\leq I\leq 7-m$. Then $M$ is CN-realizable.
@@ 4.1 Snake formation
We note that there are $J-k$ (resp. $l-I$) pairs of white edges before (resp. after) $B^{k}_{l}$ and $(J-k)+(l-I)\geq l-k-1$ because of the condition $I-J\leq 1$. Take $n,~{}m\in\mathbb{N}$ so that $n+m=k-l-1$, $n\leq J-k$, and $m\leq l-I$. (In Figure 7, we take $n=m=2$.) Apply the ladder move L7 for the black edge $B^{k}_{l}$ with the $n$ white edges $W^{k}_{k+1},W^{k+1}_{k+2},\dots$ which are just above $B^{k}_{l}$. Then, apply L3, and apply L8 for the (shortened) black edge with the $m$ white edegs $W^{l-1}_{l},W^{l-2}_{l-1},\dots$ which are below the black edge. The black edge becomes length one, and by applying L1, we obtain a W-ladder diagram. ∎
@@
Figure 8: $c$-$r$-, $c$-$rc$-, $rc$-$r$- and $rc$-$rc$-formations.
@@
Let $M$ be a matrix of the $c(k,l)$-$rc(l-1,m)$-formation. Take a B-ladder diagram $D$ of $M$ as
@@
Then, for the lower part, including the lower white edge $W^{l-1}_{l}$, apply the ladder moves in the same manner to the proof of Proposition 14 (the transformations $B$ in Figure 9) to obtain white edges. Thus, we obtain a W-ladder diagram. For the other cases, $c$-$r$-, $rc$-$r$-, $rc$-$rc$-formations, W-ladder diagrams can be obtained in the same way. ∎
@@
    Let $M$ be an $n\times n$ matrix of a $c(k,l)$-$r(l-1,m)$-formation with $l-k\geq 2$, $m-l\geq 1$. The matrix $M+M_{1}$ is also a CN-realizable matrix.
@@
    Let $M$ be an $n\times n$ matrix of a $c(k,l)$-$r(l-1,m)$-formation with $l-k\geq 2$, $m-l\geq 2$. The matrix $M+M_{1}+M_{2}$ is also a CN-realizable matrix.
We call such a formation of the matrix $M+M_{1}$ or $M+M_{1}+M_{2}$ in Proposition 16 a $c(k,l)\#r(l-1,m)$-formation (or simply a $c\#r$-formation).
@@
Figure 10: A matrix of a $c(1,4)\#r(3,7)$-formation is a CN-realizable matrix. Ignore the orange-colored broken edges if $M(2,6)=0$.
@@ 4.2 Hang-glider formation
$\left\{\begin{array}[]{ll}M(k,m+1)=M(k,l)=M(m,l)=M(m,m+1)=2.\\ M(k,j)=2\text{ when }j\leq m-1.\\ M(i,l)=2\text{ when }i\geq m+2.\\ M(k,m)=0\text{ or }2.\ M(m+1,l)=0\text{ or }2.\\ M(m-1,m+1)=0\text{ or }2.\ M(m,m+2)=0\text{ or }2.\\ M(i,j)=0\text{ otherwise}.\end{array}\right.$
@@
Let $M$ be an $n\times n$ strictly upper triangular matrix that has an $H(k,l;m)$-formation. Let $M_{1}$ (resp. $M_{2}$) be the $n\times n$ matrix such that $M_{1}(k-1,k+1)=2$ (resp. $M_{2}(l-1,l+1)=2$) and the other entries are 0. The matrices $M+M_{1}$, $M+M_{2}$, and $M+M_{1}+M_{2}$ are also CN-realizable.
@@ 4.3 Loupe formation
$\left\{\begin{array}[]{ll}M(k,j)=2\text{ when }j=k+1\text{ or }k+3\leq j\leq l.\\ M(k,k+2)=0\text{ or }2.\\ M(k+1,k+3)=0\text{ or }2.\\ M(k+2,j)=2\text{ when }k+3\leq j\leq l.\\ M(i,j)=0\text{ otherwise}.\end{array}\right.$
@@
Let $M$ be an $n\times n$ strictly upper triangular matrix that has an $L_{1}(k,l)$-formation. Let $M_{1}$ (resp. $M_{2}$) be the $n\times n$ matrix such that $M_{1}(k-1,k+1)=2$ (resp. $M_{2}(k+2,l+1)=2$) and the other entries are 0. The matrices $M+M_{1}$, $M+M_{2}$, $M+M_{1}+M_{2}$ are also CN-realizable.
@@
$\left\{\begin{array}[]{ll}M(k,j)=2\text{ when }j=k+1,k+2\text{ or}k+4\leq j\leq l.\\ M(k,k+3)=0\text{ or }2.\\ M(k+2,k+4)=0\text{ or }2.\\ M(k+3,j)=2\text{ when }k+4\leq j\leq l.\\ M(i,j)=0\text{ otherwise}.\end{array}\right.$
@@
$\left\{\begin{array}[]{ll}M(k,j)=2\text{ when }j=k+1\text{ or }k+4\leq j\leq l.\\ M(k,k+2)=0\text{ or }2.\\ M(k+1,k+4)=0\text{ or }2.\\ M(k+2,j)=2\text{ when }k+4\leq j\leq l.\\ M(k+3,j)=2\text{ when }k+4\leq j\leq l.\\ M(i,j)=0\text{ otherwise}.\end{array}\right.$
@@ 4.4 T-structure
    For each vertex $V(i,j)$ of $g$, let $m$, $n$ be the length of the vertical, horizontal paths for $V(i,j)$, respectively. Then $m+n=j-i-1$.
@@
Let $M$ be a strictly upper triangular $(0,2)$-matrix such that there is at most one pair $(i,j)$ that satisfies $j-i=k$ and $M(i,j)=2$ for each $k\geq 3$. If $M$ has a T-structure, then $M$ is a CN-realizable matrix [^2] There also exsist CN-realizable matrices with no T-structure. For example, see the matrix $M$ in Figure 5. The vertex $V(1,4)$ of the grid alignment $G(M)$ needs to have a vertical path of length two since there are no vertices on the left-hand side. The vertex $V(2,6)$ needs to have the horizontal path of length two and the vertical path of length one. Then, the undesired part consisting of the vertices $V(3,4)$, $V(2,4)$, $V(2,6)$, and $V(5,6)$ is unavoidable, which is against the condition (C2). Hence, no T-structures are allowed for $M$..
@@
The matrix $M$ shown in Figure 18 has a T-structure and we can transform a B-ladder diagram of $M$ into a W-ladder diagram by ladder moves [^3] Except for L4. based on the T-structure in the following way. Take a B-ladder diagram of $M$ so that shorter black edges are lower than the longer edges, where the length of a black edge $B^{k}_{l}$ means the value $l-k$. Apply the ladder move L1 to the black edges of length one to convert them to white edges (the transformation $A$ in Figure 18). Next, for the black edges $B^{i}_{j}$ of length 2, use the white edegs such that the corresponding vertices are on the vertical or horizontal path for $V(i,j)$ to shorten them (for example, use $W^{1}_{2}$ for $B^{1}_{3}$ in Figure 18) and apply L1 to convert them to white edges (the transformation $B$ in Figure 18). Repeat the procedure until the longest black edge is shortened and converted to white edges.
@@ 5 CN-realizable configurations
The configurations $d1$ to $d20$ shown in Figures 27 and 28 are CN-realizable configurations for a $6\times 6$ T0 upper triangular $(0,2)$-matrix.
@@ 6.3 Crossing matrix
Let $M=(M(i,j))$ be an $n\times n$ matrix. Assume that Conjecture 3 is true for $n\leq N$. Suppose that $M$ is the crossing matrix of a positive pure $n$-braid diagram $b$. Since $b$ is positive, $M$ coincides with the OU matrix of $b$ by Proposition 2, and $M$ is a non-negative integer matrix. Since $b$ is pure, $M$ is symmetric by Proposition 1. By Proposition 26, $M+M^{T}=2M$ is an even T0 matrix for $n\leq N$, and hence $M$ is a T0 symmetric matrix for $n\leq N$. Conversely, suppose that $M$ is a non-negative T0 symmetric matrix. Then, $M+M^{T}=2M$ is a non-negative even T0 (symmetric) matrix. By the assumption that Conjecture 3 is true for $n\leq N$, $2M$ is the CN matrix of some pure $n$-braid projection $B$ for $n\leq N$. The number of intersections with the strands $s_{i}$ and $s_{j}$ of $B$ is $2M(i,j)$. Let $B(i,j)$ be the number of intersections between the strands $s_{i}$ and $s_{j}$ of $B$ where $s_{i}$ crosses $s_{j}$ from left to right. (We remark that $B(i,j)+B(j,i)=2M(i,j)$.) Since $B$ is pure, $B(i,j)=B(j,i)=M(i,j)$. Let $b$ be a positive pure $n$-braid diagram whose projection is $B$. Then, for each crossing $c$ between the strands $s_{i}$ and $s_{j}$ of $b$, $s_{i}$ crosses $s_{j}$ from left to right at $c$ if and only if $s_{i}$ crosses under $s_{j}$ at $c$. Hence, the number of crossings between $s_{i}$ and $s_{j}$ of $b$ where $s_{i}$ crosses over $s_{j}$ coincides with the number of crossings between $s_{i}$ and $s_{j}$ of $b$ where $s_{i}$ crosses under $s_{j}$. Then, the $(i,j)$ entry of the OU matrix of $b$ is $B(i,j)\ (=M(i,j))$, i.e., the OU matrix of $b$ is $M$. Since $b$ is positive, the crossing matrix of $b$ is also $M$ by Proposition 2. This completes the proof of Proposition 27. ∎

### v3

@@ 2 Properties of CN matrix
for some $1\leq I\leq n-m+1$. Then $M_{2}$ is CN-realizable.
@@
Let $B_{1}$ be an $m$-braid projection such that $N(B_{1})=M_{1}$. Place I1I-1 strands on the left-hand side and nmI+1n-m-I+1 strands on the right-hand side to obtain an $n$-braid projection $B_{2}$. Then $N(B_{2})=M_{2}$. (See Figure 3.) ∎
@@
Let $M$ be a $6\times 6$ T0 strictly upper triangular $(0,2)$-matrix such that $M(i,j)=0$ when $i<I$ or j>Im1j>I-m-1 for some integers $1\leq m\leq 5$ and $1\leq I\leq 7-m$. Then $M$ is CN-realizable.
@@ 4.1 Snake formation
We note that there are JkJ-k (resp. lIl-I) pairs of white edges before (resp. after) $B^{k}_{l}$ and $(J-k)+(l-I)\geq l-k-1$ because of the condition $I-J\leq 1$. Take $n,~m\in\mathbb{N}$ so that n+m=kl1n+m=k-l-1, $n\leq J-k$, and $m\leq l-I$. (In Figure 7, we take $n=m=2$.) Apply the ladder move L7 for the black edge $B^{k}_{l}$ with the $n$ white edges $W^{k}_{k+1},W^{k+1}_{k+2},\dots$ which are just above $B^{k}_{l}$. Then, apply L3, and apply L8 for the (shortened) black edge with the $m$ white edegs $W^{l-1}_{l},W^{l-2}_{l-1},\dots$ which are below the black edge. The black edge becomes length one, and by applying L1, we obtain a W-ladder diagram. ∎
@@
Figure 8: $c$-$r$-, $c$-rcrc-, rcrc-$r$- and rcrc-rcrc-formations.
@@
Let $M$ be a matrix of the $c(k,l)$-rc(l1,m)rc(l-1,m)-formation. Take a B-ladder diagram $D$ of $M$ as
@@
Then, for the lower part, including the lower white edge $W^{l-1}_{l}$, apply the ladder moves in the same manner to the proof of Proposition 14 (the transformations $B$ in Figure 9) to obtain white edges. Thus, we obtain a W-ladder diagram. For the other cases, $c$-$r$-, rcrc-$r$-, rcrc-rcrc-formations, W-ladder diagrams can be obtained in the same way. ∎
@@
    Let $M$ be an $n\times n$ matrix of a $c(k,l)$-r(l1,m)r(l-1,m)-formation with $l-k\geq 2$, $m-l\geq 1$. The matrix $M+M_{1}$ is also a CN-realizable matrix.
@@
    Let $M$ be an $n\times n$ matrix of a $c(k,l)$-r(l1,m)r(l-1,m)-formation with $l-k\geq 2$, $m-l\geq 2$. The matrix $M+M_{1}+M_{2}$ is also a CN-realizable matrix.
We call such a formation of the matrix $M+M_{1}$ or $M+M_{1}+M_{2}$ in Proposition 16 a c(k,l)#r(l1,m)c(k,l)\#r(l-1,m)-formation (or simply a c#rc\#r-formation).
@@
Figure 10: A matrix of a c(1,4)#r(3,7)c(1,4)\#r(3,7)-formation is a CN-realizable matrix. Ignore the orange-colored broken edges if $M(2,6)=0$.
@@ 4.2 Hang-glider formation
[…]
@@
Let $M$ be an $n\times n$ strictly upper triangular matrix that has an $H(k,l;m)$-formation. Let $M_{1}$ (resp. $M_{2}$) be the $n\times n$ matrix such that $M_{1}(k-1,k+1)=2$ (resp. $M_{2}(l-1,l+1)=2$) and the other entries are 0. The matrices $M_{1}, M+M$$M_{2}$, and $M+M_{1}+M_{2}$ are also CN-realizable.
@@ 4.3 Loupe formation
[…]
@@
Let $M$ be an $n\times n$ strictly upper triangular matrix that has an $L_{1}(k,l)$-formation. Let $M_{1}$ (resp. $M_{2}$) be the $n\times n$ matrix such that $M_{1}(k-1,k+1)=2$ (resp. $M_{2}(k+2,l+1)=2$) and the other entries are 0. The matrices $M_{1}, M+M$$_{2}, M+M$$M_{1}+M_{2}$ are also CN-realizable.
@@
[…]
@@
[…]
@@ 4.4 T-structure
    For each vertex $V(i,j)$ of $g$, let $m$, $n$ be the length of the vertical, horizontal paths for $V(i,j)$, respectively. Then m+n=ji1m+n=j-i-1.
@@
Let $M$ be a strictly upper triangular $(0,2)$-matrix such that there is at most one pair $(i,j)$ that satisfies ji=kj-i=k and $M(i,j)=2$ for each $k\geq 3$. If $M$ has a T-structure, then $M$ is a CN-realizable matrix [^2] There also exsist CN-realizable matrices with no T-structure. For example, see the matrix $M$ in Figure 5. The vertex $V(1,4)$ of the grid alignment $G(M)$ needs to have a vertical path of length two since there are no vertices on the left-hand side. The vertex $V(2,6)$ needs to have the horizontal path of length two and the vertical path of length one. Then, the undesired part consisting of the vertices $V(3,4)$, $V(2,4)$, $V(2,6)$, and $V(5,6)$ is unavoidable, which is against the condition (C2). Hence, no T-structures are allowed for $M$..
@@
The matrix $M$ shown in Figure 18 has a T-structure and we can transform a B-ladder diagram of $M$ into a W-ladder diagram by ladder moves [^3] Except for L4. based on the T-structure in the following way. Take a B-ladder diagram of $M$ so that shorter black edges are lower than the longer edges, where the length of a black edge $B^{k}_{l}$ means the value lkl-k. Apply the ladder move L1 to the black edges of length one to convert them to white edges (the transformation $A$ in Figure 18). Next, for the black edges $B^{i}_{j}$ of length 2, use the white edegs such that the corresponding vertices are on the vertical or horizontal path for $V(i,j)$ to shorten them (for example, use $W^{1}_{2}$ for $B^{1}_{3}$ in Figure 18) and apply L1 to convert them to white edges (the transformation $B$ in Figure 18). Repeat the procedure until the longest black edge is shortened and converted to white edges.
@@ 5 CN-realizable configurations
The configurations $d1$ to $d21$ shown in Figures 27 and 28 are CN-realizable configurations for a $6\times 6$ T0 upper triangular $(0,2)$-matrix.
@@ 6.3 Crossing matrix
Let $M=(M(i,j))$ be an $n\times n$ matrix. Assume that Conjecture 3 is true for $n\leq N$. Suppose that $M$ is the crossing matrix of a positive pure $n$-braid diagram $b$. Since $b$ is positive, $M$ coincides with the OU matrix of $b$ by Proposition 2, and $M$ is a non-negative integer matrix. Since $b$ is pure, $M$ is symmetric by Proposition 1. By Proposition 26, $M+M^{T}=2M$ is an even T0 matrix for $n\leq N$, and hence $M$ is a T0 symmetric matrix for $n\leq N$. Conversely, suppose that $M$ is a non-negative T0 symmetric matrix. Then, $M+M^{T}=2M$ is a non-negative even T0 (symmetric) matrix. By the assumption that Conjecture 3 is true for $n\leq N$, $2M$ is the CN matrix of some pure $n$-braid projection $B$ for $n\leq N$. The number of intersections with the strands $s_{i}$ and $s_{j}$ of $B$ is $2M(i,j)$. Let $B(i,j)$ be the number of intersections between the strands $s_{i}$ and $s_{j}$ of $B$ where $s_{i}$ crosses $s_{j}$ from left to right. (We remark that $B(i,j)+B(j,i)=2M(i,j)$.) Since $B$ is pure, $B(i,j)=B(j,i)=M(i,j)$. Let $b$ be a positive pure $n$-braid diagram whose projection is $B$. Then, for each crossing $c$ between the strands $s_{i}$ and $s_{j}$ of $b$, $s_{i}$ crosses $s_{j}$ from left to right at $c$ if and only if $s_{i}$ crosses under $s_{j}$ at $c$. Hence, the number of crossings between $s_{i}$ and $s_{j}$ of $b$ where $s_{i}$ crosses over $s_{j}$ coincides with the number of crossings between $s_{i}$ and $s_{j}$ of $b$ where $s_{i}$ crosses under $s_{j}$. Then, the $(i,j)$ entry of the OU matrix of $b$ is B(i,j)(=M(i,j))B(i,j)\ (=M(i,j)), i.e., the OU matrix of $b$ is $M$. Since $b$ is positive, the crossing matrix of $b$ is also $M$ by Proposition 2. This completes the proof of Proposition 27. ∎